# Guided Flow Policy: Learning from High-Value Actions in Offline Reinforcement Learning

### v1

@@ 1 Introduction
Offline Reinforcement Learning (RL) aims to learn effective policies from static datasets without further interaction with the environment wiering2012reinforcement; ernst2005tree. This paradigm is essential in domains such as robotics and logistics, where online exploration can be unsafe or costly. However, standard off-policy algorithms such as DDPG lillicrap2015continuous and SAC haarnoja2018soft, which are successful in online RL, tend to underperform in offline settings since the RL agent cannot interact with the environment. The main challenge is extrapolation error, corresponding to the inability to properly evaluate out-of-distribution actions wu2019behavior; fujimoto2019off; kumar2019stabilizing; kumar2020conservative.
@@
Two main lines of work have been proposed to address this challenge. The first one focuses on learning a critic without querying the values of actions outside the dataset kostrikov2021offline; nair2020awac. The second one, known as the Behavior-Regularized Actor–Critic (BRAC) family [^1] The abbreviation BRAC refers to the family of methods, not solely the BRAC paper wu2019behavior., mitigates these errors by forcing the learned policy to stay ”close” to the unknown behavior policy that generated the dataset fujimoto2021minimalist; tarasov2023revisiting; jaques2019way; laroche2019safe; wu2019behavior. The key idea is that out-of-distribution state–action pairs are especially vulnerable to Q-value overestimation, while staying near the empirical distribution reduces extrapolation errors. Minimalist variants achieve this by simply adding a behavior cloning (BC) loss to the policy and/or value updates with respect to dataset actions fujimoto2021minimalist; tarasov2023revisiting. Although this approach improves stability, it also raises a trade-off: regularizing too strictly to a potentially suboptimal dataset action may restrict the policy from exploiting higher-reward actions contained in the dataset.
@@
Until recently, most offline RL algorithms were based on Gaussian policies, with limited modeling capacities. Recent development of flow and diffusion-based expressive models lipman2022flow; ho2020denoising; song2020score, led to new RL algorithms to capture complex and multimodal action distributions chi2023diffusion; janner2022planning; wang2022diffusion; zhang2025energy. However, they come at the risk of high computational overhead: iterative sampling slows inference, and directly optimizing the values of output actions would result in unstable backpropagation through time (BPTT). Among recent approaches to address these challenges, park2025flow proposed a BRAC method with a flow-matching BC model distilled into a one-step policy that also optimizes the critic, enabling expressive policy learning while avoiding BPTT and iterative sampling at inference. Despite these advances, a central limitation remains: the flow-based BC component, similar to standard BC, does not incorporate reward information.
@@
We propose Guided Flow Policy (GFP), a dual-policy BRAC framework with a bidirectional guidance mechanism between a multi-step flow-matching policy, termed Value-aware Behavior Cloning (VaBC), and a distilled one-step actor. VaBC acts as a distributional regularizer for the actor, encouraging it to remain within the support of the behavior policy. VaBC is trained via a weighted-BC mechanism, close to peng2019advantage; zhang2025energy, but leveraging the actor and the critic to prioritize cloning high-value actions from the dataset. Unlike previous BRAC approaches, in which the BC regularization indiscriminately treats all state-action pairs fujimoto2021minimalist; park2025flow, VaBC is a value-aware regularizer integrated into a BRAC approach. In turn, the actor optimizes the critic while being distilled toward VaBC, allowing it to align with the dataset’s high-value actions in a given state while maximizing expected returns. Fig. 1 illustrates the GFP framework and Tab. 1 shows how it differs in the regularization mechanisms compared to other BRAC methods.
@@
Our contributions are threefold: (i) we introduce Guided Flow Policy, a simple yet effective BRAC method that integrates value-awareness in the regularization term via a jointly trained weighted flow BC policy, thereby regularizing the actor with the most promising transitions of the dataset; (ii) we extensively evaluate GFP on 144 tasks from standard offline RL benchmarks, showing strong performances with substantial gains on suboptimal datasets and challenging tasks compared to prior works; (iii) we re-assess two previous state-of-the-art offline RL algorithms on these benchmarks, highlighting the critical role of hyperparameter choices and subtle implementation details, aligned in the spirit with the retrospective analysis provided in tarasov2023revisiting.
@@
Figure 1: Overview of the Guided Flow Policy framework. GFP consists of three main components: (i) in yellow, VaBC, a multi-step flow policy $\pi_{\omega}$ trained via weighted BC using the guidance term $g_{\eta}$, (ii) in green, a one-step actor $\pi_{\theta}$ distilled from the flow policy, and (iii) in gray, a critic $Q_{\phi}$ guiding action evaluation. $\pi_{\omega}$ regularizes the actor toward high-value actions from the dataset; in turn, the actor shapes the flow and optimizes the critic following the actor–critic approach. The different components of the figure are introduced throughout the paper. Each drawing represents the probability distribution of actions $a\in\mathcal{A}$ of a policy, in a current state $s$, except for the gray ones, where it is the value of actions $a\in\mathcal{A}$ in state $s$, according to the critic.
@@
Regularization target Value-aware regularization Expressive variant TD3+BC fujimoto2021minimalist ReBRAC tarasov2023revisiting $\hookrightarrow$ dataset actions Diffusion-QL wang2022diffusion - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - FQL park2025flow $\hookrightarrow$ learned behavior cloning policy - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - GFP (ours) $\hookrightarrow$ learned value-aware behavior cloning policy
@@ 2 Background
Actor-critic framework in RL. RL problems are typically formalized as a Markov Decision Process (MDP) sutton1998reinforcement; konda1999actor, defined by the tuple $(\mathcal{S},\mathcal{A},p,r,\rho,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p$ the transition dynamics, $r$ the reward function, $\rho$ the initial state distribution, and $\gamma\in[0,1)$ the discount factor. The behavior of the agent is governed by a policy $\pi$, mapping states to probability distributions over actions. The objective is to maximize the expected discounted return $\mathbb{E}_{a_{t}\sim\pi(\cdot\mid s_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, i.e., the expected cumulative reward when following $\pi$ in the MDP. In actor-critic approaches, the policy $\pi$, also referred to as the actor, is trained jointly with a critic $Q$, which approximates the state–action value function. The $Q$-function is defined as $Q^{\pi}(s,a)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t},a_{t}\right)\mid s_{0}=s,a_{0}=a\right]$, estimating the expected return after taking action $a$ in state $s$ and subsequently following $\pi$.
@@
Minimalist approaches in offline RL. In offline RL, the agent learns exclusively from a static dataset $\mathcal{D}$, consisting of transitions $(s,a,r,s^{\prime})$ generated by an unknown behavior policy. In a given state $s$, the distribution of actions of such a behavior policy is illustrated on the left of Fig. 1. This introduces a key challenge compared to the online setting: the distributional shift wiering2012reinforcement; kumar2019stabilizing; konda1999actor. Indeed, since the learned policy $\pi_{\theta}$ may select actions outside the dataset’s support, value estimates for such out-of-distribution actions can be inaccurate kumar2020conservative; fujimoto2021minimalist. The BRAC approach addresses this issue by constraining the policy $\pi_{\theta}$ to remain close to the behavior policy jaques2019way; kumar2019stabilizing. As emphasized in fujimoto2021minimalist, a simple and effective choice is to add a BC term directly into the actor objective. Incorporating this into the actor–critic framework, the actor loss in Eq. 1 becomes:
@@
Behavior cloning with flow matching. Flow Matching (FM) lipman2022flow is a generative modeling framework that learns a continuous-time transformation, or flow, which maps a simple base distribution (in this work, a standard Gaussian) to a target data distribution. This transformation is defined through a family of intermediate, time-dependent distributions governed by an ordinary differential equation (ODE).
@@
This flow, conditioned on the state $s$, maps noise samples $z\sim\mathcal{N}(0,I_{d})$ into actions distributed according to $\pi_{\omega}(\cdot\mid s)$. While sophisticated conditioning strategies can help enhance expressiveness (e.g., classifier-free guidance ho2022classifier), we adopt in this work the simplest variant of conditional flow matching holderrieth2025introduction. We further employ the optimal transport variant of FM, which uses linear interpolation with uniformly sampled time points lipman2022flow. For $(s,a)\sim\mathcal{D}$, $\epsilon\sim\mathcal{N}(0,I_{d})$, and $t\sim\mathcal{U}([0,1])$, we define the interpolated point $a_{t}=(1-t)\epsilon+ta$, whose target velocity is $a-\epsilon$. The velocity field $v_{\theta}$ is then trained by least-squares regression toward this reference, yielding the conditional flow-matching BC loss holderrieth2025introduction:
@@
Flow policy for offline RL. Following the idea of Diffusion Q-Learning wang2022diffusion, a straightforward way to train a flow policy for offline RL is to replace the BC term in the actor loss (Eq. 3) with the flow-matching BC loss (Eq. 5). However, the iterative sampling procedure makes training expensive, due to recursive backpropagation through time (BPTT) in the actor loss, and also results in a slow inference at test time. To mitigate these limitations, park2025flow suggests distilling the iterative flow-matching BC policy into a one-step policy that also maximizes the critic.
@@ 3 Guided Flow Policy
We now detail the GFP algorithm that builds on top of fujimoto2021minimalist; park2025flow. GFP integrates a Value-aware Behavior Cloning (VaBC) flow policy with a distilled one-step actor through bidirectional guidance. VaBC leverages the actor and the critic to selectively clone high-value dataset actions, providing more targeted regularization than in standard BRAC approaches. The distilled actor, in turn, maximizes the critic while avoiding BPTT and iterative sampling. GFP is composed of three main components: the critic $Q_{\phi}$, the actor $\pi_{\theta}$, and the VaBC policy $\pi_{\omega}$. The complete algorithm is presented in Algo. 1, and the approach is illustrated in Fig. 1.
@@
Figure 2: Comparison of behavior cloning under different levels of guidance. Left: Prior work (e.g., FQL, park2025flow) uses no filtering, indiscriminately imitating all state-action pairs. Right: In contrast, our method introduces a temperature-controlled guidance mechanism, as shown in Eq. 10, resulting in VaBC. At high temperatures, the guidance is weak, so the actor is influenced by many candidate actions. At moderate temperatures, the filtering becomes sharper, giving more weight to higher-value actions while still keeping enough regularization and exploration. At low temperatures, the filtering is very selective, concentrating almost exclusively on the highest-value actions according to the critic. However, excessive concentration at very low temperatures may allow the actor to escape the dataset’s action distribution, as shown on the right in green, leading to critic overestimation and out-of-distribution issues. Importantly, VaBC cannot escape the dataset’s action distribution even at very low temperatures, since it trains exclusively on in-distribution state-action pairs. The dashed blue contours in the final yellow drawings (first row) illustrate this constraint.
@@
$\mathcal{L^{\mathcal{A}}}(\theta)=\mathbb{E}_{s\sim\mathcal{D},z\sim\mathcal{N}(0,I_{d})}\Big[-\lambda Q_{\phi}\left(s,\mu_{\theta}(s,z)\right)+\alpha\|\mu_{\theta}(s,z)-\mu_{\bar{\omega}}(s,z)\|_{2}^{2}\Big].$ (8)
The normalization term $\lambda=\frac{1}{\frac{1}{N}\sum|Q_{\phi}(s,a)|}$ is based on the average absolute Q-value, estimated over mini-batches rather than over the entire dataset fujimoto2021minimalist.
@@
The key contribution of GFP is to add value-awareness in the behavior regularization component of a BRAC framework, effectively combining the two predominant policy extraction methods studied by park2024value: weighted-BC and behavior regularized policy gradients, further presented in Sec. 5. One could employ an advantage weighted term, similar to AWR peng2019advantage, using the actor to compute a baseline:
@@
As recommended by peng2019advantage, in that case, a clipping term should be added for stability. Instead, we propose the guiding function $g_{\eta}$, which corresponds to a soft-max between $Q_{\phi}(s,a)$ and $Q_{\phi}(s,\mu_{\theta}(s,z))$. In appendix, Sec. B.3, Tab. 7 reports results when using $g^{\text{AWR}}_{\eta}$ instead of $g_{\eta}$, over 65 tasks.

### v2

@@ 1 Introduction
Offline Reinforcement Learning (RL) aims to learn effective policies from static datasets without further interaction with the environment (S. Lange, 2012; Ernst et al., 2005). This paradigm is essential in domains such as robotics and logistics, where online exploration can be unsafe or costly. However, standard off-policy algorithms such as DDPG (Lillicrap et al., 2015) and SAC (Haarnoja et al., 2018), which are successful in online RL, tend to underperform in offline settings since the RL agent cannot interact with the environment. The main challenge is extrapolation error, corresponding to the inability to properly evaluate out-of-distribution actions (Wu et al., 2019; Fujimoto et al., 2019; Kumar et al., 2019; 2020).
@@
Two main lines of work have been proposed to address this challenge. The first one focuses on learning a critic without querying the values of actions outside the dataset (Kostrikov et al., 2021; Nair et al., 2020). The second one, known as the Behavior-Regularized Actor–Critic (BRAC) family [^1] The abbreviation BRAC refers to the family of methods, not solely the BRAC paper (Wu et al., 2019)., mitigates these errors by forcing the learned policy to stay ”close” to the unknown behavior policy that generated the dataset (Fujimoto and Gu, 2021; Tarasov et al., 2023; Jaques et al., 2019; Laroche et al., 2019; Wu et al., 2019). The key idea is that out-of-distribution state–action pairs are especially vulnerable to Q-value overestimation, while staying near the empirical distribution reduces extrapolation errors. Minimalist variants achieve this by simply adding a behavior cloning (BC) loss to the policy and/or value updates with respect to dataset actions (Fujimoto and Gu, 2021; Tarasov et al., 2023). Although this approach improves stability, it also raises a trade-off: regularizing too strictly to a potentially suboptimal dataset action may restrict the policy from exploiting higher-reward actions contained in the dataset.
@@
Until recently, most offline RL algorithms were based on Gaussian policies, with limited modeling capacities. Recent development of flow and diffusion-based expressive models (Lipman et al., 2022; Ho et al., 2020; Song et al., 2020), led to new RL algorithms to capture complex and multimodal action distributions (Chi et al., 2023; Janner et al., 2022; Wang et al., 2022; Zhang et al., 2025). However, they come at the risk of high computational overhead: iterative sampling slows inference, and directly optimizing the values of output actions would result in unstable backpropagation through time (BPTT). Among recent approaches to address these challenges, Park et al. (2025) proposed a BRAC method with a flow-matching BC model distilled into a one-step policy that also optimizes the critic, enabling expressive policy learning while avoiding BPTT and iterative sampling at inference. Despite these advances, a central limitation remains: the flow-based BC component, similar to standard BC, does not incorporate reward information.
@@
We propose Guided Flow Policy (GFP), a dual-policy BRAC framework with a bidirectional guidance mechanism between a multi-step flow-matching policy, termed Value-aware Behavior Cloning (VaBC), and a distilled one-step actor. VaBC acts as a distributional regularizer for the actor, encouraging it to remain within the support of the behavior policy. VaBC is trained via a weighted-BC mechanism, close to Peng et al. (2019); Zhang et al. (2025), but leveraging the actor and the critic to prioritize cloning high-value actions from the dataset. Unlike previous BRAC approaches, in which the BC regularization indiscriminately treats all state-action pairs (Fujimoto and Gu, 2021; Park et al., 2025), VaBC is a value-aware regularizer integrated into a BRAC approach. In turn, the actor optimizes the critic while being distilled toward VaBC, allowing it to align with the dataset’s high-value actions in a given state while maximizing expected returns. Fig. 1 illustrates the GFP framework and Tab. 1 shows how it differs in the regularization mechanisms compared to other BRAC methods.
@@
Our contributions are threefold: (i) we introduce Guided Flow Policy, a simple yet effective BRAC method that integrates value-awareness in the regularization term via a jointly trained weighted flow BC policy, thereby regularizing the actor with the most promising transitions of the dataset; (ii) we extensively evaluate GFP on 144 tasks from standard offline RL benchmarks, showing strong performances with substantial gains on suboptimal datasets and challenging tasks compared to prior works; (iii) we re-assess two previous state-of-the-art offline RL algorithms on these benchmarks, highlighting the critical role of hyperparameter choices and subtle implementation details, aligned in the spirit with the retrospective analysis provided in Tarasov et al. (2023).
@@
Figure 1: Overview of the Guided Flow Policy framework. GFP consists of three main components: (i) in yellow, VaBC, a multi-step flow policy $\pi_{\omega}$ trained via weighted BC using the guidance term $g_{\eta}$, (ii) in green, a one-step actor $\pi_{\theta}$ distilled from the flow policy, and (iii) in gray, a critic $Q_{\phi}$ guiding action evaluation. $\pi_{\omega}$ regularizes the actor toward high-value actions from the dataset $\mathcal{D}$; in turn, the actor shapes the flow and optimizes the critic following the actor–critic approach. The different components of the figure are introduced throughout the paper. Each drawing represents the probability distribution of actions $a\in\mathcal{A}$ of a policy, in a current state $s$, except for the gray ones, where it is the value of actions $a\in\mathcal{A}$ in state $s$, according to the critic.
@@
Regularization target Value-aware regularization Expressive variant TD3+BC (Fujimoto and Gu, 2021) ReBRAC (Tarasov et al., 2023) $\hookrightarrow$ dataset actions Diffusion-QL (Wang et al., 2022) - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - FQL (Park et al., 2025) $\hookrightarrow$ learned behavior cloning policy - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - GFP (ours) $\hookrightarrow$ learned value-aware behavior cloning policy
@@ 2 Background
Actor-critic framework in RL. RL problems are typically formalized as a Markov Decision Process (MDP) (Sutton et al., 1998; Konda and Tsitsiklis, 1999), defined by the tuple $(\mathcal{S},\mathcal{A},p,r,\rho,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p$ the transition dynamics, $r$ the reward function, $\rho$ the initial state distribution, and $\gamma\in[0,1)$ the discount factor. The behavior of the agent is governed by a policy $\pi$, mapping states to probability distributions over actions. The objective is to maximize the expected discounted return $\mathbb{E}_{a_{t}\sim\pi(\cdot\mid s_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, i.e., the expected cumulative reward when following $\pi$ in the MDP. In actor-critic approaches, the policy $\pi$, also referred to as the actor, is trained jointly with a critic $Q$, which approximates the state–action value function. The $Q$-function is defined as $Q^{\pi}(s,a)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}\left[\sum_{t=0}^{\infty}\gamma^{t}r\left(s_{t},a_{t}\right)\mid s_{0}=s,a_{0}=a\right]$, estimating the expected return after taking action $a$ in state $s$ and subsequently following $\pi$.
@@
Minimalist approaches in offline RL. In offline RL, the agent learns exclusively from a static dataset $\mathcal{D}$, consisting of transitions $(s,a,r,s^{\prime})$ generated by an unknown behavior policy. In a given state $s$, the distribution of actions of such a behavior policy is illustrated on the left of Fig. 1. This introduces a key challenge compared to the online setting: the distributional shift (S. Lange, 2012; Kumar et al., 2019; Konda and Tsitsiklis, 1999). Indeed, since the learned policy $\pi_{\theta}$ may select actions outside the dataset’s support, value estimates for such out-of-distribution actions can be inaccurate (Kumar et al., 2020; Fujimoto and Gu, 2021). The BRAC approach addresses this issue by constraining the policy $\pi_{\theta}$ to remain close to the behavior policy (Jaques et al., 2019; Kumar et al., 2019). As emphasized in Fujimoto and Gu (2021), a simple and effective choice is to add a BC term directly into the actor objective. Incorporating this into the actor–critic framework, the actor loss in Eq. 1 becomes:
@@
Behavior cloning with flow matching. Flow Matching (FM) (Lipman et al., 2022) is a generative modeling framework that learns a continuous-time transformation, or flow, which maps a simple base distribution (in this work, a standard Gaussian) to a target data distribution. This transformation is defined through a family of intermediate, time-dependent distributions governed by an ordinary differential equation (ODE).
@@
This flow, conditioned on the state $s$, maps noise samples $z\sim\mathcal{N}(0,I_{d})$ into actions distributed according to $\pi_{\omega}(\cdot\mid s)$. While sophisticated conditioning strategies can help enhance expressiveness (e.g., classifier-free guidance (Ho and Salimans, 2022)), we adopt in this work the simplest variant of conditional flow matching (Holderrieth and Erives, 2025). We further employ the optimal transport variant of FM, which uses linear interpolation with uniformly sampled time points (Lipman et al., 2022). For $(s,a)\sim\mathcal{D}$, $\epsilon\sim\mathcal{N}(0,I_{d})$, and $t\sim\mathcal{U}([0,1])$, we define the interpolated point $a_{t}=(1-t)\epsilon+ta$, whose target velocity is $a-\epsilon$. The velocity field $v_{\theta}$ is then trained by least-squares regression toward this reference, yielding the conditional flow-matching BC loss (Holderrieth and Erives, 2025):
@@
Flow policy for offline RL. Following the idea of Diffusion Q-Learning by Wang et al. (2022), a straightforward way to train a flow policy for offline RL is to replace the BC term in the actor loss (Eq. 3) with the flow-matching BC loss (Eq. 5). However, the iterative sampling procedure makes training expensive, due to recursive backpropagation through time (BPTT) in the actor loss, and also results in a slow inference at test time. To mitigate these limitations, Park et al. (2025) suggests distilling the iterative flow-matching BC policy into a one-step policy that also maximizes the critic.
@@ 3 Guided Flow Policy
We now detail the GFP algorithm that builds on top of Fujimoto and Gu (2021); Park et al. (2025). GFP integrates a Value-aware Behavior Cloning (VaBC) flow policy with a distilled one-step actor through bidirectional guidance. VaBC leverages the actor and the critic to selectively clone high-value dataset actions, providing more targeted regularization than in standard BRAC approaches. The distilled actor, in turn, maximizes the critic while avoiding BPTT and iterative sampling. GFP is composed of three main components: the critic $Q_{\phi}$, the actor $\pi_{\theta}$, and the VaBC policy $\pi_{\omega}$. The complete algorithm is presented in Algo. 1, and the approach is illustrated in Fig. 1.
@@
Figure 2: Comparison of behavior cloning under different levels of guidance. Left: Prior work (e.g., FQL by Park et al. (2025)) uses no filtering, indiscriminately imitating all state-action pairs. Right: In contrast, our method introduces a temperature-controlled guidance mechanism, as shown in Eq. 10, resulting in VaBC. At high temperatures, the guidance is weak, so the actor is influenced by many candidate actions. At moderate temperatures, the filtering becomes sharper, giving more weight to higher-value actions while still keeping enough regularization and exploration. At low temperatures, the filtering is very selective, concentrating almost exclusively on the highest-value actions according to the critic. However, excessive concentration at very low temperatures may allow the actor to escape the dataset’s action distribution, as shown on the right in green, leading to critic overestimation and out-of-distribution issues. Importantly, VaBC cannot escape the dataset’s action distribution even at very low temperatures, since it trains exclusively on in-distribution state-action pairs. The dashed blue contours in the final yellow drawings (first row) illustrate this constraint.
@@
$\mathcal{L^{\mathcal{A}}}(\theta)=\mathbb{E}_{s\sim\mathcal{D},z\sim\mathcal{N}(0,I_{d})}\Big[-\lambda Q_{\phi}\left(s,\mu_{\theta}(s,z)\right)+\alpha\|\mu_{\theta}(s,z)-\mu_{\omega}(s,z)\|_{2}^{2}\Big].$ (8)
The normalization term $\lambda=\frac{1}{\frac{1}{N}\sum|Q_{\phi}(s,a)|}$ is based on the average absolute Q-value, estimated over mini-batches rather than over the entire dataset (Fujimoto and Gu, 2021).
@@
The key contribution of GFP is to add value-awareness in the behavior regularization component of a BRAC framework, effectively combining the two predominant policy extraction methods studied by Park et al. (2024b): weighted-BC and behavior regularized policy gradients, further presented in Sec. 5. One could employ an advantage weighted term, similar to AWR (Peng et al., 2019), using the actor to compute a baseline:
@@
As recommended by Peng et al. (2019), in that case, a clipping term should be added for stability. Instead, we propose the guiding function $g_{\eta}$, which corresponds to a soft-max between $Q_{\phi}(s,a)$ and $Q_{\phi}(s,\mu_{\theta}(s,z))$. In appendix, Sec. B.3, Tab. 7 reports results when using $g^{\text{AWR}}_{\eta}$ instead of $g_{\eta}$, over 65 tasks.